OpenStudy (anonymous):

Classify the conic section and write its equation in standard form. 1) y2 - 8y - 8x = -24

OpenStudy (anonymous):

That's a parabola, but I forgot what is it standard form. Let me google it.

OpenStudy (anonymous):

parabola

OpenStudy (anonymous):

Complete the square on the LHS with: y^2 - 8y + 24 = 8x. Then divide by 8.

OpenStudy (anonymous):

is it x = - yea

OpenStudy (anonymous):

I think it's \(\LARGE \frac{(y - 4)^2}{8} + 1 = x \) but I don't recall if the coefficient for the (y - 4) has to be 1 or not.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yea - not sure - when the subject is introduced it uses y^2 = 4 a x

OpenStudy (anonymous):

Indeed, so I guess we need to expand it out without and simplifying then squaring again?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It should be x = ay^2, so yeah, guess we have to work it out.

OpenStudy (anonymous):

i'm not sure - the literature often says that y = ax^2 + bx + c is the standard form so is x = ay^2 + by + c considered to be std form as well - so your first suggestion could be right

OpenStudy (anonymous):

Yeah, guess it's a problem with ambiguity. Do you have answer choices, @lindawalters?